how many grams of calcium ions are in 0.5000g of calcium hydroxide
1
Expert's answer
2012-11-09T06:29:12-0500
M(Ca(OH)2) = 40 + 17*2 = 74 g/mol
w(Ca2+) = 40/74 * 100% = 54%
m(Ca2+) = 0.5000 * 0.54 = 0.27 g